if 10.0g of AI and 5.00g of HCI are initially present,find the:



a.)limiting reactant



b.)excess reactant



c.)mass of excess reactant that will be consumed in the reaction

1
Expert's answer
2022-03-25T12:44:02-0400

2Al + 6HCl----------> 2AlCl3 + 3H2

a) Limiting reactant is HCl

B) Excess reactant is Al

C) Moles of Al = 10/27= 0.37037 moles

Moles of HCl= 5/36.5 = 0.13697 moles

Excess moles of Al= (0.37037-0.13697)= 0.2334 moles

Mass of excess Al= 0.2334×27 =6.3018g